OpenStudy (anonymous):

Find the probability of rolling a sum of five with a single toss of two fair six-sided dice.

OpenStudy (anonymous):

@Hero

hero (hero):

Okay...

OpenStudy (anonymous):

??

OpenStudy (anonymous):

is it 1/6

OpenStudy (anonymous):

what so i do count all the dice that have 6

hero (hero):

There are 36 possible pairs, and of those, there are four pairs that sum to five

hero (hero):

So the probability of rolling a sum of five with two six-sided dice is \[\frac{4}{36} = \frac{1}{9}\]

OpenStudy (anonymous):

so its1/9 what about this one Find the probability of getting one head and two tails on a toss of three fair coins

hero (hero):

The same logic applies. Find the total possible pairs, then find how many of those have one head and two tails.

hero (hero):

It's only three coins so you can probably do it with pencil and paper.

hero (hero):

here's how I figured out the previous one: Notice how I circled those that add to five.

OpenStudy (anonymous):

would it be 1/8

hero (hero):

HHH HHT HTH `HTT` THH `THT` `TTH` TTT

hero (hero):

There are eight possible pairs, and of those, three of them have meet our requirements.

OpenStudy (anonymous):

so 1/3 would be the correct answer

hero (hero):

\[\text{probability} = \frac{\text{no of occurrences}}{\text{total possible outcomes}}\]

hero (hero):

There are eight possible outcomes, of those only three of the expected event occurs.

OpenStudy (anonymous):

ohh wow its 3/8 lol can you help me on this one too If we roll two dice 180 times, how many times can we expect the dice to show a sum of five? 5 20 45 36

hero (hero):

I can only help you so much before you'd be expected to start figuring these out on your own.

OpenStudy (anonymous):

is it the same for this one the way you did it before

hero (hero):

Here, solve this proportion: \[\frac{4}{36} = \frac{x}{180}|\] It's not the same approach.

OpenStudy (anonymous):

x=20

hero (hero):

Good

OpenStudy (anonymous):

ok i think i get it now :)

hero (hero):

You'd have to prove to me that you get any of this.